### v1

# Xing Huang(a), Eva Kopfer(b), Pierre Monmarché(c), Panpan Ren(d) a)

Abstract
In this paper, we derive exponential ergodicity in relative entropy for general kinetic SDEs under a partially dissipative condition. It covers non-equilibrium situations where the forces are not of gradient type and the invariant measure does not have an explicit density, extending previous results set in the equilibrium case. The key argument is to establish the hypercontractivity of the associated semigroup, which follows from its hyperboundedness and its $L^{2}$-exponential ergodicity. Moreover, we obtain exponential ergodicity in the $L^{2}$-Wasserstein distance by combining Talagrand’s inequality with a log-Harnack inequality. These results are further extended to the McKean-Vlasov setting and to the associated mean-field interacting particle systems, with convergence rates that are uniform in the number of particles in the latter case, under small nonlinear perturbations.
AMS subject Classification:  60H10, 60K35, 82C22.
Keywords: Hypercontractivity, exponential ergodicity, Non-equilibrium kinetic SDEs, relative entropy, $L^{2}$-Wasserstein distance, McKean-Vlasov SDEs, mean field interacting particle system
1 Introduction
The (kinetic) Langevin equation is a central model of statistical physics and molecular dynamics. It describes the motion of a classical particle subject to external forces and coupled with a Brownian heat bath. In the so-called equilibrium case, the forces derive from a potential and the Langevin equation reads
(1.3) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=-\gamma Y_{t}\text{\rm{d}}t-\nabla V(X_{t})\text{\rm{d}}t+\sqrt{2\gamma\beta^{-1}}\text{\rm{d}}W_{t}\,,\end{array}\right.$
where $\gamma>0$ is the friction parameter, $\beta^{-1}$ is the inverse temperature, $V$ is the external (or confining) potential and $W$ is a multi-dimensional Brownian motion. Here, $X\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}^{d}$ respectively stands for the position and velocity of the particle. This equation has been extensively studied and we refer e.g. to [32] and references within for general considerations. Specifically, in the present work, we are interested in the long-time relaxation to a statistical steady state for the process. Under mild conditions on $V$, (1.3) is known to be ergodic with respect to the Gibbs measure $\mu$ with density proportional to $\text{\rm{e}}^{-\beta H}$ with $H(x,y)=V(x)+|y|^{2}/2$ (assuming $e^{-\beta H}\in L^{1}$). Exponential ergodicity under suitable conditions was first established in [56, 38], using Harris theorem.
The first exponential convergence result stated in terms of relative entropy has been obtained in the seminal work [58] of Villani. It introduced the notion of hypocoercivity which has been developped since then in various directions, see [18, 44, 1, 10, 6, 8] and references therein on this general topic. More specifically, [58, Theorem 39] gives an exponential decay in relative entropy along (1.3) provided $V\in C^{2}(\mathbb{R}^{d})$ and has a bounded Hessian, and $\mu$ satisfies a so-called log-Sobolev inequality (see (3.3) below) with respect to the Dirichlet form $\mathscr{E}(f,g)=\mu(\nabla f\cdot\nabla g)$, where $\nabla$ is the gradient on $\mathbb{R}^{2d}$.
Apart from its intrinsic role in statistical physics, a key advantage of the relative entropy with respect to total variation, $V$-norms or $L^{2}$ norms is that it scales well with dimension, in particular for mean-field systems and thus, by extension, is suitable for the non-linear limits of the latter; moreover it also behaves well with respect to discretization schemes. This explains that obtaining quantitative convergence rates for the Langevin equation in relative entropy, in particular in mean-field cases, is still a very active topic [9, 13, 14, 15, 19, 22, 28, 36, 47, 49, 51, 54, 57] (motivated in particular by the mean-field analysis of high-dimensionl algorithms).
However, the result in [58, Theorem 39], along with nearly all subsequent works, relies on having an explicit representation of the invariant probability measure $\mu$. An exception arises in the so-called uniformly dissipative case, which is a very restrictive situation where the drift of the SDE induces a deterministic contraction along the stochastic flow [5, 40, 42]. In particular, there are many tools to establish log-Sobolev inequalities [3], but most of them require an explicit expression, or are restricted to the uniformly-dissipative case as the Bakry-Émery criterion (which in its classical form is written for explicit invariant measures but in fact apply more generally [42]).
Situations where $\mu$ has no explicit form are known as non-equilibrium models [23]. This is for instance the case if the conservative force $-\nabla V$ is replaced by a non-gradient force field (as in [21, 31, 30]), or if the temperature $\beta^{-1}$ is not constant among different coordinates (as in [7, 33]). In contrast to the equilibrium setting, the position and velocity are not independent under $\mu$ in these non-equilibrium situations [46]. There are also non-equilibrium examples in mean-field situations, for instance when the interaction between particles is not symmetric as in some plasma models [11, 45], mean-field adversarial games [35, 62] or other situations with multi-type particles [37].
Although recent works such as [50, 17, 43] have established exponential ergodicity in $L^{2}$-distances under potentially restrictive conditions, proving exponential convergence in relative entropy for non-equilibrium kinetic stochastic differential equations (SDEs) with only partially dissipative drift has remained an open problem for a long time. This work addresses that problem and further extends the analysis to nonlinear McKean–Vlasov equations and the corresponding interacting particle systems.
The remainder of this work is organized as follows. In Section 2 we present our main results, Theorems 2.3, 2.6 and 2.7. More specifically, Theorem 2.3 establishes exponential ergodicity in both relative entropy and the $L^{2}$-Wasserstein distance for a broad class of kinetic SDEs under a partially dissipative condition. Theorems 2.6 and 2.7 extend these results to the McKean–Vlasov setting and to the associated mean-field interacting particle systems, with convergence rates that are uniform in the number of particles in the latter case, under small nonlinear perturbations. Section 3 is devoted to the proofs of these results.
2 Main results
In the following we present the main results of this article.
Let $\mathscr{P}(\mathbb{R}^{2d})$ be all the probability measures on $\mathbb{R}^{2d}$ equipped with the weak topology. For $p\geq 1$, let
$\mathscr{P}_{p}(\mathbb{R}^{2d}):=\big{\{}\mu\in\mathscr{P}(\mathbb{R}^{2d}):\|\mu\|_{p}:=\mu(|\cdot|^{p})^{\frac{1}{p}}<\infty\big{\}},$
which is a Polish space under the $L^{p}$-Wasserstein distance
$\mathbb{W}_{p}(\mu,\nu)=\inf_{\pi\in\mathscr{C}(\mu,\nu)}\bigg{(}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{p}\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg{)}^{\frac{1}{p}},\ \ \mu,\nu\in\mathscr{P}_{p}(\mathbb{R}^{2d}),$
where $\mathscr{C}(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$.
For $\nu,\mu\in\mathscr{P}(\mathbb{R}^{2d})$, we write $\mathrm{Ent}(\nu|\mu)$ the relative entropy of $\nu$ with respect to $\mu$. The law of a random variable $Z$ is denoted by $\mathscr{L}_{Z}$.
2.1 Exponential Ergodicity for classical kinetic SDEs
Our first main result provides exponential ergodicity for classical kinetic SDEs. Let $\sigma\in\mathbb{R}^{d}\otimes\mathbb{R}^{n}$, $b:\mathbb{R}^{2d}\rightarrow\mathbb{R}^{d}$ be measurable and locally bounded and $\{W_{t}\}_{t\geq 0}$ be an $n$-dimensional standard Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$.
We make the following assumption.
  1. (B)
    There exists a constant $K_{b}>0$ such that for all $z,\bar{z}\in\mathbb{R}^{2d}$,
    (2.1) $|b(z)-b(\bar{z})|\leq K_{b}|z-\bar{z}|.$
    Moreover, there exist constants $\theta,r,R>0,$ and $r_{0}\in(-1,1)$ such that for any $x,y,\bar{x},\bar{y}\in\mathbb{R}^{d}$ satisfying $|x-\bar{x}|^{2}+|y-\bar{y}|^{2}\geq R^{2}$,
    $\langle r^{2}(x-\bar{x})+rr_{0}(y-\bar{y}),y-\bar{y}\rangle$
    (2.2) $+\langle(y-\bar{y})+rr_{0}(x-\bar{x}),b(x,y)-b(\bar{x},\bar{y})\rangle$
    $\leq-\theta(|x-\bar{x}|^{2}+|y-\bar{y}|^{2}).$
Remark 2.1.
Condition (B) is equivalent to [43, Assumption  1], which can be seen by noting that $r_{0}\in(-1,1)$ and considering the matrix
$A=\left(\begin{array}[]{cc}r^{2}I_{d\times d}&rr_{0}I_{d\times d}\\ rr_{0}I_{d\times d}&I_{d\times d}\\ \end{array}\right)\,.$
Example 2.2.
As detailed in [43, Example 1], for any $\gamma>0$, there exists $\kappa>0$ such that (B) is satisfied for $b(x,y)=-x-\gamma y+F(x,y)$ provided $|\nabla F|\leqslant\kappa$ outside a compact set.
We consider the kinetic SDE
(2.5) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=b(X_{t},Y_{t})\text{\rm{d}}t+\sigma\text{\rm{d}}W_{t}.\end{array}\right.$
Note that (2.1) implies that (2.5) is well-posed. We denote by $(X_{t}^{z},Y_{t}^{z})$ the solution to (2.5) with initial value $z\in\mathbb{R}^{2d}$ and by $(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})$ the solution starting from $\mu_{0}\in\mathscr{P}$. We further denote by $P_{t}$ the associated semigroup.
Theorem 2.3.
Assume (B) and $\sigma\sigma^{\ast}$ is invertible. Then (2.5) has a unique invariant probability measure $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$ and there exists a constant $t_{1}>0$ such that $P_{t_{1}}^{\ast}$, the adjoint operator of $P_{t_{1}}$ in $L^{2}(\bar{\mu})$, is hypercontractive, in the sense that
(2.6) $\|P_{t_{1}}^{\ast}\|_{L^{\frac{4}{3}}(\bar{\mu})\rightarrow L^{2}(\bar{\mu})}=1.$
Consequently, there exist constants $c,\tilde{c},\lambda>0$ such that
(2.7) $\mathrm{Ent}(\mathscr{L}_{(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})}|\bar{\mu})\leq c\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\bar{\mu}),~{}t\geq 0,\mu_{0}\in\mathscr{P}(\mathbb{R}^{2d}),$
and
(2.8) $\mathbb{W}_{2}(\mathscr{L}_{(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})},\bar{\mu})\leq\tilde{c}\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Remark 2.4.
In the case $n=d$, $\sigma=\sigma_{0}I_{d\times d}$ for some positive constant $\sigma_{0}$, $L^{2}$-exponential ergodicity for (2.5) has been derived in [43]. In fact, the result in [43] can be easily extended to the general case of $\sigma\sigma^{\ast}$ being an invertible $d\times d$ matrix, as we shall see in Lemma 3.3 below.
Remark 2.5.
In the uniform dissipative case, [59] uses hypercontractivity to derive the exponential ergodicity in relative entropy. Let us remark that in the uniform dissipative case, the hypercontractivity can be derived by Wang’s Haranck inequality with power while in the partially dissipative case, Wang’s Harnack inequality can only derive hyperboundedness, see [28]. Recently, [43] derives the $L^{2}$-exponential ergodicity under the partially dissipative condition. Note that exponential ergodicity in relative entropy is strictly stronger than $L^{2}$-exponential ergodicity, see for instance [59, Proposition 2.3]. Together with hyperboundedness, $L^{2}$-exponential ergodicity implies exponential ergodicity in relative entropy, as we shall recall in Lemma 3.1.
2.2 Exponential ergodicity for McKean-Vlasov SDEs
Our second main result generalizes exponential ergodicity towards distribution dependent SDEs. Let $b:\mathbb{R}^{2d}\times\mathscr{P}_{2}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}$, $\sigma:\mathscr{P}_{2}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ be measurable and bounded on bounded sets. We assume the following.
  1. (C)
    There exist constants $K_{b}>0,K_{I}>0$ such that for all $z,\bar{z}\in\mathbb{R}^{2d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$,
    (2.9) $|b(z,\gamma)-b(\bar{z},\tilde{\gamma})|+\|\sigma(\gamma)-\sigma(\tilde{\gamma})\|_{HS}\leq K_{b}|z-\bar{z}|+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma}).$
    Moreover, there exist constants $\theta,r,R>0,$ and $r_{0}\in(-1,1)$ such that for any $x,y,\bar{x},\bar{y}\in\mathbb{R}^{d}$ satisfying $|x-\bar{x}|^{2}+|y-\bar{y}|^{2}\geq R^{2}$ and $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$,
    $\langle r^{2}(x-\bar{x})+rr_{0}(y-\bar{y}),y-\bar{y}\rangle$
    (2.10) $+\langle(y-\bar{y})+rr_{0}(x-\bar{x}),b(x,y,\mu)-b(\bar{x},\bar{y},\mu)\rangle$
    $\leq-\theta(|x-\bar{x}|^{2}+|y-\bar{y}|^{2}).$
    In addition, there exist constants $0<\delta_{2}\leq\delta_{1}$ such that
    (2.11) $\delta_{2}\leq\sigma\sigma^{\ast}\leq\delta_{1}.$
We consider the McKean-Vlasov SDE
(2.14) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=b(X_{t},Y_{t},\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}t+\sigma(\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}W_{t}.\end{array}\right.$
First we note that under (2.9), (2.14) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$. We use $P_{t}^{\ast}\mu_{0}$ to denote the distribution of the solution to (2.14) with initial distribution $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$.
Theorem 2.6.
Assume (C). There exist $K_{*},c,\tilde{c},\lambda>0$, depending on $d$ and on the constants in (C) except on $K_{I}$, such that, provided $K_{I}\leqslant K_{*}$, then (2.14) has a unique stationary solution $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$,
(2.15) $\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\bar{\mu})\leq c\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),$
and
(2.16) $\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\bar{\mu})\leq\tilde{c}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\bar{\mu}),~{}t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
2.3 Exponential approximate ergodicity for mean field interacting particle system
Our third main result provides an approximate exponential ergodicity for mean field interacting kinetic particle systems of the form
(2.19) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}^{i,N}=Y_{t}^{i,N}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}^{i,N}=b\left(X_{t}^{i,N},Y_{t}^{i,N},\frac{1}{N}\sum_{i=1}^{N}\delta_{(X_{t}^{i,N},Y_{t}^{i,N})}\right)\text{\rm{d}}t+\sigma\text{\rm{d}}W_{t}^{i},\end{array}\right.$
where $(W_{t}^{i})_{i\geq 1}$ are independent $n$-dimensional Brownian motions. In this case, we are not able to consider the case where $\sigma$ depends on the empirical measure of the particles, since the Harnack inequality for multiplicative kinetic SDEs is still open. The corresponding McKean-Vlasov SDE is
(2.22) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=b(X_{t},Y_{t},\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}t+\sigma\text{\rm{d}}W_{t}.\end{array}\right.$
Let us denote by $(P_{t}^{N})^{\ast}\nu^{N}$ the distribution of the solution to (2.19) with initial distribution $\nu^{N}\in\mathscr{P}((\mathbb{R}^{2d})^{N})$.
Theorem 2.7.
Assume (C). There exist $K_{*},c,\tilde{c},\lambda>0$, depending on $d$ and on the constants in (C) except on $K_{I}$, such that, provided $K_{I}\leqslant K_{*}$ then, for all $N\geqslant 1$, (2.19) has a unique invariant probability measure $\bar{\mu}^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N})$,
(2.23) $\mathbb{W}_{2}((P_{t}^{N})^{\ast}\nu^{N},\bar{\mu}^{N})^{2}\leq c\text{\rm{e}}^{-2\lambda t}\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{N})^{2}+cNR_{d}(N),\ \ t\geq 0,\nu^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N}),$
and
(2.24) $\mathrm{Ent}((P_{t}^{N})^{\ast}\nu^{N}|\bar{\mu}^{N})\leq\tilde{c}\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{N})^{2}+cNR_{d}(N),~{}t\geq 1,\nu^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N})$
where
$\begin{split}R_{d}(N)=\begin{cases}N^{-\frac{1}{2}},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}d<2,\\ N^{-\frac{1}{2}}\log(1+N),~{}~{}~{}~{}~{}~{}d=2,\\ N^{-\frac{2}{d}},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}d>2.\end{cases}\end{split}$
Remark 2.8.
Compared with the existing results on the exponential ergodicity in $L^{2}$-Wasserstein distance and relative entropy for second-order McKean-Vlasov SDEs/mean field interacting particle system uniform in the number of particles under partially dissipative condition (such as [27, 13] ; see also [25] for the case with non-degenerate additive noise), the results in Theorems 2.6 and 2.7 are the first where the drift term is not of gradient form and the invariant probability measure for the associated mean field interacting particle system has no explicit density.
Remark 2.9.
When $K_{I}$ is small enough, (2.22) admits a unique stationary probability distribution $\bar{\mu}$ thanks to Theorem 2.6. Along the proof of Theorem 2.7, we establish that $\mathcal{W}_{2}(\bar{\mu}^{\otimes N},\bar{\mu}^{N})^{2}\leqslant CNR_{N}$ for some constant $C>0$. Using triangular inequality, this shows that (2.23) also holds if we replace $\bar{\mu}^{N}$ by $\bar{\mu}^{\otimes N}$. Then, if particles are indistinguishable (i.e. if $\nu^{N}$ is invariant by permutation of the particles’ labels), writing $\nu_{t}^{k,N}$ the marginal law of $k\leqslant N$ particles under the law $(P_{t}^{N})^{\ast}\nu^{N}$ and using the scaling properties of $\mathbb{W}_{2}$, we get
$\mathbb{W}_{2}(\nu_{t}^{k,N},\bar{\mu}^{\otimes k})^{2}\leq c\text{\rm{e}}^{-2\lambda t}\mathbb{W}_{2}(\nu_{0}^{k,N},\bar{\mu}^{\otimes k})^{2}+ckR_{d}(N),\ \ t\geq 0.$
Example 2.10.
As in Example 2.2, for any $\gamma>0$, there exists $\kappa>0$ such that the conditions in Theorem 2.6 and Theorem 2.7 are satisfied for
$b(z,\mu)=-x-\gamma y+F_{1}(x,y)+\int_{\mathbb{R}^{2d}}F_{2}(z,\bar{z})\mu(\text{\rm{d}}\bar{z}),\ \ z=(x,y)\in\mathbb{R}^{2d}$
with $F_{1},F_{2}:\mathbb{R}^{2d}\times\mathbb{R}^{2d}\rightarrow\mathbb{R}^{d}$ provided $|\nabla F_{1}|\leqslant\kappa$, $|\nabla_{1}F_{2}|\leqslant\kappa$ outside a compact set and $\|\nabla_{2}F_{2}\|_{\infty}$ is small enough.
3 Proof of the Main Theorems
Before we go on and prove our main results, we shall mention the following key result for our analysis. In order to do so, consider a conservative diffusion process on $\mathbb{R}^{m}$ with generator $L=\frac{1}{2}\mathrm{tr}(a\nabla^{2})+\langle b,\nabla\rangle,$ and associated semigroup $P_{t}$. Let $\mu$ be an invariant probability measure of $(P_{t})_{t\geq 0}$. By [59, Proposition 2.3], the exponential convergence for $\mu(P_{t}f\log P_{t}f)$ can be derived from hypercontractivity of $P_{t}$, i.e. $\|P_{t_{0}}\|_{L^{2}(\mu)\rightarrow L^{4}(\mu)}\leq 1$ for some $t_{0}>0$. Moreover, in that case, the entropy convergence rate depends only on $t_{0}$. On the other hand, [4, Step (d) in the proof of Theorem 1.1] states that hypercontractivity of $P_{t}$ can be deduced by the exponential convergence of $\mu((P_{t}f)^{2})$ together with hyperboundedness $\|P_{\tilde{t}_{0}}\|_{L^{2}(\mu)\rightarrow L^{4}(\mu)}\leq C$ for some constants $\tilde{t}_{0}>0,C>0$. Building on [59, Proposition 2.3] and [4, Step (d) in Proof of Theorem 1.1], we summarize the following key result on exponential ergodicity in relative entropy.
Lemma 3.1.
Let $\bar{\mu}$ be the invariant probability measure of $(P_{t})_{t\geq 0}$, and assume that $(P_{t})_{t\geq 0}$ is hyperbounded, i.e., $\|P_{t_{0}}\|_{p\rightarrow q}:=\|P_{t_{0}}\|_{L^{p}(\bar{\mu})\rightarrow L^{q}(\bar{\mu})}<\infty$ for some $t_{0}>0$ and $1<p<q<\infty$. Assume moreover that $L^{2}$-exponential ergodicity holds, i.e.
$\|P_{t}f-\bar{\mu}(f)\|_{L^{2}(\bar{\mu})}\leq c\text{\rm{e}}^{-\lambda t}\|f-\bar{\mu}(f)\|_{L^{2}(\bar{\mu})},\ \ t\geq 0$
for some constants $c,\lambda>0$. Then hypercontractivity holds, i.e.
$\|P_{\tilde{t}_{0}}\|_{p\rightarrow q}\leq 1$
for some constant $\tilde{t}_{0}>0$ depending only on $c,\lambda,q,p,t_{0}$ and $\|P_{t_{0}}\|_{p\rightarrow q}$. Consequently, exponential ergodicity in relative entropy holds:
$\bar{\mu}(P_{t}f\log P_{t}f)\leq\tilde{c}\text{\rm{e}}^{-\tilde{\lambda}t}\bar{\mu}(f\log f),\ \ \bar{\mu}(f)=1,f>0,t\geq 0$
for some constants $\tilde{c},\tilde{\lambda}>0$ depending only on $\tilde{t_{0}},q,p$.
Remark 3.2.
An introduction to hypercontractivity can be found in [2].
3.1 Proof of Theorem 2.3
In this section, we prove exponential ergodicity in relative entropy and $L^{2}$-Wasserstein distance for classical kinetic SDEs.
We first present a lemma on the $L^{2}$-exponential ergodicity for (2.5). The proof largely follows the approach in [43, Theorem 2], where the technique of hypocoercivity plays a central role.
Lemma 3.3.
Assume (2.1) and $\sigma\sigma^{\ast}$ is invertible. If in addition (2.5) has a unique invariant probability measure $\bar{\mu}$ such that the Poincaré inequality
$\bar{\mu}(f^{2})-\bar{\mu}(f)^{2}\leq C_{\mathrm{PI}}\bar{\mu}(|\nabla f|^{2}),\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
holds for some constant $C_{\mathrm{PI}}$, then there exist constants $c_{0}>0,\lambda_{0}>0$ depending on $C_{PI}$, $K_{b}$ and the minimum eigenvalue of $\sigma\sigma^{\ast}$ such that
$\|P_{t}f-\bar{\mu}(f)\|_{L^{2}(\bar{\mu})}\leq c_{0}\text{\rm{e}}^{-\lambda_{0}t}\|f-\bar{\mu}(f)\|_{L^{2}(\bar{\mu})},\ \ t\geq 0,\ f\in L^{2}(\mu).$
Proof.
Let
$M=2(2K_{b}+1)^{2}+(2K_{b}+1).$
Let $\delta_{1}$ be the minimum eigenvalue of $\sigma\sigma^{\ast}$ and $\varepsilon=\frac{\delta_{1}}{M+\frac{1}{2}}$. By an approximation technique, we can assume that $b\in C^{1}$ with $\|\nabla b\|_{\infty}<K_{b}$. For $f\in C_{0}^{\infty}(\mathbb{R}^{2d})$, let $f_{t}=P_{t}f-\bar{\mu}(f)$. Then it follows from the fact
(3.1) $L(g^{2})=2gLg+|\sigma^{\ast}\nabla_{y}g|^{2},\ \ \bar{\mu}(L(g^{2}))=0,\ \ g\in C_{0}^{\infty}(\mathbb{R}^{2d})$
that
$\frac{\text{\rm{d}}\|f_{t}\|_{L^{2}(\bar{\mu})}^{2}}{\text{\rm{d}}t}=2\bar{\mu}(f_{t}Lf_{t})=-\bar{\mu}(|\sigma^{\ast}\nabla_{y}f_{t}|^{2}).$
Let $\alpha(t)=1-\text{\rm{e}}^{-t/3}$
$G_{t}=\varepsilon\left(\begin{array}[]{cc}\alpha^{3}(t)I_{d\times d}&-\alpha^{2}(t)I_{d\times d}\\ -\alpha^{2}(t)I_{d\times d}&\alpha(t)I_{d\times d}\\ \end{array}\right),$
and denote the Jacobi matrix of the drift as
$J=\left(\begin{array}[]{cc}0_{d\times d}&\nabla_{x}b\\ I_{d\times d}&\nabla_{y}b\\ \end{array}\right).$
Then it is easy to see that
$\frac{\text{\rm{d}}G_{t}}{\text{\rm{d}}t}=\varepsilon\left(\begin{array}[]{cc}3\alpha^{2}(t)\alpha^{\prime}(t)I_{d\times d}&-2\alpha(t)\alpha^{\prime}(t)I_{d\times d}\\ -2\alpha(t)\alpha^{\prime}(t)I_{d\times d}&\alpha^{\prime}(t)I_{d\times d}\\ \end{array}\right)$
and
$2G_{t}J$ $=2\varepsilon\left(\begin{array}[]{cc}\alpha^{3}(t)I_{d\times d}&-\alpha^{2}(t)I_{d\times d}\\ -\alpha^{2}(t)I_{d\times d}&\alpha(t)I_{d\times d}\\ \end{array}\right)\left(\begin{array}[]{cc}0_{d\times d}&\nabla_{x}b\\ I_{d\times d}&\nabla_{y}b\\ \end{array}\right)$
$=2\varepsilon\left(\begin{array}[]{cc}-\alpha^{2}(t)I_{d\times d}&\alpha^{3}(t)\nabla_{x}b-\alpha^{2}(t)\nabla_{y}b\\ \alpha(t)I_{d\times d}&-\alpha^{2}(t)\nabla_{x}b+\alpha(t)\nabla_{y}b\\ \end{array}\right).$
Moreover, for $z=(x,y)\in\mathbb{R}^{2d}$, it holds
$\langle G_{t}z,z\rangle=\alpha(t)(\alpha^{2}(t)|x|^{2}-2\alpha(t)\langle x,y\rangle+|y|^{2})\leq 2\varepsilon|z|^{2}.$
Observing
$\nabla Lf_{t}=L\nabla f_{t}+J\nabla f_{t},$
and in view of (3.1),
$\bar{\mu}(\langle G_{t}L\nabla f_{t},\nabla f_{t}\rangle)=\bar{\mu}(\langle L\sqrt{G_{t}}\nabla f_{t},\sqrt{G_{t}}\nabla f_{t}\rangle)=-\bar{\mu}(|\sigma^{\ast}\nabla_{y}\sqrt{G_{t}}\nabla f_{t}|^{2})\leq 0,$
we conclude that
$\frac{\text{\rm{d}}\bar{\mu}(\langle G_{t}\nabla f_{t},\nabla f_{t}\rangle)}{\text{\rm{d}}t}$
$=\bar{\mu}(\langle\frac{\text{\rm{d}}G_{t}}{\text{\rm{d}}t}\nabla f_{t},\nabla f_{t}\rangle)+2\bar{\mu}(\langle G_{t}\nabla Lf_{t},\nabla f_{t}\rangle)$
$=\bar{\mu}(\langle\frac{\text{\rm{d}}G_{t}}{\text{\rm{d}}t}\nabla f_{t},\nabla f_{t}\rangle)+2\bar{\mu}(\langle G_{t}L\nabla f_{t},\nabla f_{t}\rangle)+2\bar{\mu}(\langle G_{t}J\nabla f_{t},\nabla f_{t}\rangle)$
$\leq\bar{\mu}(\langle\frac{\text{\rm{d}}G_{t}}{\text{\rm{d}}t}\nabla f_{t},\nabla f_{t}\rangle)+2\bar{\mu}(\langle G_{t}J\nabla f_{t},\nabla f_{t}\rangle).$
Set
$\mathscr{N}_{t}=\|f_{t}\|_{L^{2}(\bar{\mu})}^{2}+\bar{\mu}(\langle G_{t}\nabla f_{t},\nabla f_{t}\rangle).$
Then we have
$\frac{\text{\rm{d}}\mathscr{N}_{t}}{\text{\rm{d}}t}\leq\bar{\mu}\left(R_{t}\nabla f_{t},\nabla f_{t}\right)$
where
$R_{t}$ $=-\left(\begin{array}[]{cc}0_{d\times d}&0_{d\times d}\\ 0_{d\times d}&\sigma\sigma^{\ast}\end{array}\right)+\frac{\text{\rm{d}}G_{t}}{\text{\rm{d}}t}+2G_{t}J$
$=\left(\begin{array}[]{cc}\varepsilon(3\alpha^{\prime}(t)-2)\alpha^{2}(t)I_{d\times d}&\varepsilon(2\alpha^{3}(t)\nabla_{x}b-2\alpha^{2}(t)\nabla_{y}b-2\alpha(t)\alpha^{\prime}(t))\\ \varepsilon 2\alpha(t)(1-\alpha^{\prime}(t))&-\sigma\sigma^{\ast}+\varepsilon\alpha^{\prime}(t)-\varepsilon\alpha^{2}(t)\nabla_{x}b+\varepsilon\alpha(t)\nabla_{y}b\end{array}\right).$
It is not difficult to see from $\alpha^{\prime}(t)\in(0,1/3]$ and $\alpha\in[0,1)$ that
$\langle R_{t}z,z\rangle$ $=\varepsilon(3\alpha^{\prime}(t)-2)\alpha^{2}(t)|x|^{2}+\langle[\varepsilon(2\alpha^{3}(t)\nabla_{x}b-2\alpha^{2}(t)\nabla_{y}b-2\alpha(t)\alpha^{\prime}(t))]y,x\rangle$
$+\langle[\varepsilon 2\alpha(t)(1-\alpha^{\prime}(t))x,y\rangle+\langle[-\sigma\sigma^{\ast}+\varepsilon\alpha^{\prime}(t)-\varepsilon\alpha^{2}(t)\nabla_{x}b+\varepsilon\alpha(t)\nabla_{y}b]y,y\rangle$
$\leq-\varepsilon\alpha^{2}(t)|x|^{2}-\langle\sigma\sigma^{\ast}y,y\rangle$
$+\varepsilon 2\alpha(t)|x|(\|\nabla_{x}b\|_{\infty}+\|\nabla_{y}b\|_{\infty}+1)|y|+\varepsilon(\|\nabla_{x}b\|_{\infty}+\|\nabla_{y}b\|_{\infty}+1)|y|^{2}$
$\leq-\frac{\varepsilon\alpha^{2}(t)}{2}|x|^{2}-\delta_{1}|y|^{2}$
$+\varepsilon\{2(\|\nabla_{x}b\|_{\infty}+\|\nabla_{y}b\|_{\infty}+1)^{2}+(\|\nabla_{x}b\|_{\infty}+\|\nabla_{y}b\|_{\infty}+1)\}|y|^{2}$
$\leq-\frac{\varepsilon\alpha^{2}(t)}{2}|x|^{2}+(\varepsilon M-\delta_{1})|y|^{2}.$
Therefore, we get
$\frac{\text{\rm{d}}\mathscr{N}_{t}}{\text{\rm{d}}t}\leq-\frac{\varepsilon\alpha^{2}(t)}{2}\bar{\mu}(|\nabla f_{t}|^{2}).$
Observe that it follows from Poincaré’s inequality that
$\bar{\mu}(|\nabla f_{t}|^{2})\geq\frac{1}{C_{P_{I}}}\bar{\mu}(|f_{t}|^{2})=\frac{1}{C_{P_{I}}}[\mathscr{N}_{t}-\bar{\mu}(\langle G_{t}\nabla f_{t},\nabla f_{t}\rangle)]\geq\frac{1}{C_{P_{I}}}[\mathscr{N}_{t}-2\varepsilon\bar{\mu}(|\nabla f_{t}|^{2})],$
which implies
$\bar{\mu}(|\nabla f_{t}|^{2})\geq\frac{1}{C_{P_{I}}+2\varepsilon}\mathscr{N}_{t}.$
Consequently, we derive
$\frac{\text{\rm{d}}\mathscr{N}_{t}}{\text{\rm{d}}t}\leq-\frac{\varepsilon\alpha^{2}(t)}{2C_{P_{I}}+4\varepsilon}\mathscr{N}_{t}.$
Gronwall’s inequality and $\mathscr{N}_{0}=\|f_{0}\|_{L^{2}(\bar{\mu})}^{2}$ yields
$\|f_{t}\|_{L^{2}(\bar{\mu})}^{2}\leq\mathscr{N}_{t}\leq\exp\left\{-\frac{\varepsilon}{2C_{P_{I}}+4\varepsilon}\int_{0}^{t}\alpha(s)^{2}\text{\rm{d}}s\right\}\|f_{0}\|_{L^{2}(\bar{\mu})}^{2}.$
The proof is completed by noting that for $t\geq 1$,
$\int_{0}^{t}\alpha^{2}(s)\text{\rm{d}}s\geq\int_{1}^{t}(1-\text{\rm{e}}^{-1/3})^{2}\text{\rm{d}}s=(1-\text{\rm{e}}^{-1/3})^{2}(t-1).$
Proof of Theorem 2.3.
(Step 1) The existence and uniqueness of an invariant probability measure is known, see for instance [28, Proof of Theorem 3.3]. By [28, Theorem 3.4], there exist constants $t_{0}>0$, $C_{0}>0$ such that
(3.2) $\|P_{t_{0}}\|_{L^{2}(\bar{\mu})\rightarrow L^{4}(\bar{\mu})}\leq C_{0}.$
Moreover, it follows from [28, Theorem 3.5] that the log-Sobolev inequality holds:
(3.3) $\bar{\mu}(f\log f)\leq C_{1}\bar{\mu}(|\nabla\sqrt{f}|^{2}),\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),f>0,\bar{\mu}(f)=1,$
for some $C_{1}>0$. This immediately implies the Poincaré inequality,
$\bar{\mu}(f^{2})-\bar{\mu}(f)^{2}\leq C_{1}\bar{\mu}(|\nabla f|^{2}),\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}).$
According to Lemma 3.3, the $L^{2}$-exponential ergodicity holds, i.e. there exist $c_{0},\lambda_{0}>0$ such that
(3.4) $\|P_{t}f-\bar{\mu}(f)\|_{L^{2}(\bar{\mu})}\leq c_{0}\text{\rm{e}}^{-\lambda_{0}t}\|f-\bar{\mu}(f)\|_{L^{2}(\bar{\mu})},\ \ t\geq 0,\ f\in L^{2}(\mu).$
Hence, (3.4) together with (3.2) and Lemma 3.1 implies hypercontractivity for large enough $t_{1}\geq t_{0}$,
(3.5) $\|P_{t_{1}}\|_{L^{2}(\bar{\mu})\rightarrow L^{4}(\bar{\mu})}=1.$
(Step 2) Next, we prove the hypercontractivity of $P_{t_{1}}^{\ast}$. For any $f\in L^{\frac{4}{3}}(\bar{\mu})$, set
$f_{n}=(-n\vee f)\wedge n,\ \ n\geq 1.$
Then $f_{n}\in L^{2}(\bar{\mu})$, which together with the fact that $P_{t_{1}}^{\ast}$ is the adjoint operator of $P_{t}$ in $L^{2}(\bar{\mu})$ implies that
$\|P_{t_{1}}^{\ast}|f_{n}|\|_{L^{2}(\bar{\mu})}=\sup_{g\in L^{2}(\bar{\mu})}\int_{\mathbb{R}^{2d}}(P_{t_{1}}^{\ast}|f_{n}|)g\text{\rm{d}}\bar{\mu}$ $=\sup_{g\in L^{2}(\bar{\mu})}\int_{\mathbb{R}^{2d}}|f_{n}|P_{t_{1}}g\text{\rm{d}}\bar{\mu}$
$\leq\sup_{g\in L^{2}(\bar{\mu})}\|P_{t_{1}}g\|_{L^{4}(\bar{\mu})}\|f_{n}\|_{L^{\frac{4}{3}}(\bar{\mu})}$
$\leq\|P_{t_{1}}\|_{L^{2}(\bar{\mu})\rightarrow L^{4}(\bar{\mu})}\|f_{n}\|_{L^{\frac{4}{3}}(\bar{\mu})}=\|f_{n}\|_{L^{\frac{4}{3}}(\bar{\mu})}.$
By Fatou’s lemma and dominated convergence theorem, we derive (2.6), which combined with Lemma 3.1 implies (2.7).
(Step 3) By (2.1), it is standard to derive
(3.6) $\sup_{t\in[0,1]}\mathbb{W}_{2}(\mathscr{L}_{(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})},\bar{\mu})\leq c_{0}\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
for some constant $c_{0}>0$. By [53], a log-Harnack inequality holds:
$P_{1}\log f(x)\leq\log P_{1}f(y)+c_{1}|x-y|^{2},$
for some $c_{1}>0$. By [61, Proposition 1.4.4(3)] this implies
(3.7) $\mathrm{Ent}(\mathscr{L}_{(X_{1}^{\mu_{0}},Y_{1}^{\mu_{0}})},\mathscr{L}_{(X_{1}^{\nu_{0}},Y_{1}^{\nu_{0}})})\leq c_{1}\mathbb{W}_{2}(\mu_{0},\nu_{0})^{2}.$
Moreover, by [52], (3.3) implies the Talagrand inequality:
(3.8) $\mathbb{W}_{2}(\nu,\bar{\mu})^{2}\leq 4C_{1}\mathrm{Ent}(\nu|\bar{\mu}),\ \ \nu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
For simplicity, denote $\mu_{t}=\mathscr{L}_{(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})}$. Finally, for any $t\geq 1$, by (3.8), (2.7) and (3.7), we have
$\mathbb{W}_{2}(\mathscr{L}_{(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})},\bar{\mu})^{2}$ $\leq 4C_{1}\mathrm{Ent}(\mathscr{L}_{(X_{t}^{\mu_{0}},Y_{t}^{\mu_{0}})}|\bar{\mu})$
$\leq 4C_{1}c\text{\rm{e}}^{-2\lambda(t-1)}\mathrm{Ent}(\mathscr{L}_{(X_{1}^{\mu_{0}},Y_{1}^{\mu_{0}})}|\bar{\mu})$
$\leq 4C_{1}c\text{\rm{e}}^{-2\lambda(t-1)}c_{1}\mathbb{W}_{2}(\mu_{0},\bar{\mu})^{2}.$
Combining this with (3.6) completes the proof of (2.8), hence of Theorem 2.3. ∎
3.2 Proof of Theorem 2.6
In this section, we will prove exponential ergodicity in relative entropy and $\mathbb{W}_{2}$ for the McKean-Vlasov SDE (2.14). Recall that $P_{t}^{\ast}\mu_{0}$ denotes the distribution of the solution to (2.14) with initial distribution $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$.
For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, consider the time-homogeneous decoupled SDEs
(3.11) $\left\{\begin{array}[]{ll}\text{\rm{d}}\tilde{X}_{t}^{\mu}=\tilde{Y}_{t}^{\mu}\text{\rm{d}}t,\\ \text{\rm{d}}\tilde{Y}_{t}^{\mu}=b(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu},\mu)\text{\rm{d}}t+\sigma(\mu)\text{\rm{d}}W_{t}.\end{array}\right.$
Under (C), (3.11) is well-posed and has a unique invariant probability measure $\Phi(\mu)\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, see for instance [28, Theorem 3.3]. Let $\tilde{P}_{t}^{\mu}$ be the associated semigroup to (3.11). For any $M>0$, let
$\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}):=\left\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}):\|\mu\|_{2}^{2}\leq M\right\}.$
We first give a log-Sobolev inequality for $\Phi(\mu)$ uniform over $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$, which is a corollary of [28, Theorem 3.5(2)].
Lemma 3.4.
Assume (C). Then for any $M>0$, the log-Sobolev inequality holds:
$\Phi(\mu)(f\log f)\leq C_{\mathrm{LS}}(M)\Phi(\mu)(|\nabla\sqrt{f}|^{2}),$
(3.12) $\qquad\quad f\in C_{0}^{\infty}(\mathbb{R}^{2d}),f>0,\Phi(\mu)(f)=1,\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}),$
where $C_{\mathrm{LS}}(M)$ is a constant depending on $M$, $d$ and the constants in (C).
Proof.
By [28, Theorem 3.5(2)], it is sufficient to prove that for any $R>0$ and $M>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $M$ such that
(3.13) $C_{1}\leq\rho_{\mu}(x)\leq C_{2},\ \ x\in B_{R},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$
for $\rho_{\mu}:=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x}$. For $\lambda>0$, let
$g_{\lambda}(1,z)=\text{\rm{e}}^{-\frac{|z|^{2}}{2\lambda}},\ \ z\in\mathbb{R}^{2d}.$
For any $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$, let $\theta_{t}^{\mu}(z)=(\theta_{t}^{(1),\mu}(z),\theta_{t}^{(2),\mu}(z))$ solve
$\left\{\begin{array}[]{ll}\text{\rm{d}}\theta_{t}^{(1),\mu}(z)=\theta_{t}^{(2),\mu}(z)\text{\rm{d}}t,\\ \text{\rm{d}}\theta_{t}^{(2),\mu}(z)=b(\theta_{t}^{\mu}(z),\mu)\text{\rm{d}}t,\ \ \theta_{0}^{\mu}(z)=z\in\mathbb{R}^{2d}.\end{array}\right.$
Since
(3.14) $|b(z,\mu)|\leq K_{b}|z|+K_{I}\sqrt{M}+|b(0,\delta_{0})|,\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}),z\in\mathbb{R}^{2d},$
we conclude that
(3.15) $|\theta_{1}^{\mu}(z)|\leq\bar{C}|z|+C_{M},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}),z\in\mathbb{R}^{2d}$
for some constant $C_{M}>0$ depending on $M$. Moreover, by (3.14), (C) and [12, Theorem 1.1(1)], there exist $C_{0}>1$ and $\lambda_{0}>0$ depending on $M$ such that the heat kernel $\tilde{p}_{t}^{\mu}(z,\tilde{z})$ associated to (3.11) satisfies
$C_{0}^{-1}g_{\lambda_{0}^{-1}}(1,\tilde{z}-\theta_{1}^{\mu}(z))\leq\tilde{p}_{1}^{\mu}(z,\tilde{z})\leq C_{0}g_{\lambda_{0}}(1,\tilde{z}-\theta_{1}^{\mu}(z)),$
(3.16) $\ \ z,\tilde{z}\in\mathbb{R}^{2d},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
Moreover, it follows from [28, (3.12)] that $(\Phi(\mu))_{\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})}$ is tight so that there exists a constant $R_{0}>0$ such that
$\int_{B_{R_{0}}}\rho_{\mu}(z)\text{\rm{d}}z>\frac{1}{2},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
Note that
$\rho_{\mu}(\tilde{z})=\int_{\mathbb{R}^{2d}}\tilde{p}_{1}^{\mu}(z,\tilde{z})\rho_{\mu}(z)\text{\rm{d}}z,$
and for any $z,\tilde{z}\in\mathbb{R}^{2d},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$,
$0\leq|\mathbb{T}_{1}^{-1}(\tilde{z}-\theta_{1}^{\mu}(z))|^{2}=|\tilde{z}-\theta_{1}^{\mu}(z)|^{2}\leq 2|\tilde{z}|+2|\theta_{1}^{\mu}(z)|^{2}\leq\tilde{C}_{N}+C_{3}|\tilde{z}|^{2}+C_{3}|z|^{2}.$
We derive from (3.2) that $\rho_{\mu}(\tilde{z})\leq C_{0}$ and
(3.17) $\rho_{\mu}(\tilde{z})\geq\int_{B_{R_{0}}}\tilde{p}_{1}^{\mu}(z,\tilde{z})\rho_{\mu}(z)\text{\rm{d}}z\geq C_{1}(M,R),\ \ \tilde{z}\in B_{R},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
Hence, we obtain (3.13) and the proof is complete. ∎
With the above lemma in hand, we now present a result on the uniform $L^{2}$-exponential ergodicity in $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$ for any $M>0$.
Lemma 3.5.
Assume (C). Then for any $M>0$, there exist constants $c_{0}>0,\lambda_{0}>0$ depending on $M$, $d$ and the constants in (C) such that
$\|\tilde{P}_{t}^{\mu}f-\Phi(\mu)(f)\|_{L^{2}(\Phi(\mu))}\leq c_{0}\text{\rm{e}}^{-\lambda_{0}t}\|f-\Phi(\mu)(f)\|_{L^{2}(\Phi(\mu))},\ \ t\geq 0,\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
Proof.
Firstly, it holds
(3.18) $|b(z,\mu)-b(z,\bar{\mu})|\leq K_{b}|z-\bar{z}|,\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Second, (3.4) implies the Poincaré inequality
$\Phi(\mu)(f^{2})-\Phi(\mu)(f)^{2}\leq C_{\mathrm{PI}}(M)\Phi(\mu)(|\nabla f|^{2}),\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
This combined with (3.18), (2.11) and Lemma 3.3 completes the proof. ∎
Proof of Theorem 2.6.
(Step 1: Existence and Uniqueness of invariant probability measure)
(i) Thanks to [28, Theorem 3.3], there exists a constant $\eta_{0}>0$ such that when $K_{I}<\eta_{0}$, we can find a constant $M>0$ such that $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$ is fixed by the map $\Phi$. By [28, Theorem 3.4], there exist constants $t_{0}>0$ and $C_{0}>0$ such that
(3.19) $\|\tilde{P}_{t_{0}}^{\mu}\|_{L^{2}(\Phi(\mu))\rightarrow L^{4}(\Phi(\mu))}\leq C_{0},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
Moreover, $t_{0}$ and $C_{0}$ can be chosen independently from $K_{I}\in[0,\eta_{0})$, since if the condition (C) is satisfied for some $K_{I}<\eta_{0}$ then it is also satisfied wit $K_{I}$ replaced by $\eta_{0}$. This together with Lemmas 3.5 and 3.1 implies that for large enough $t_{1}\geq t_{0}$ (again, independent from $K_{I}$),
(3.20) $\|\tilde{P}_{t_{1}}^{\mu}\|_{L^{2}(\Phi(\mu))\rightarrow L^{4}(\Phi(\mu))}=1,\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
Let $(\tilde{P}_{t}^{\mu})^{\ast}$ stand for the adjoint operator of $\tilde{P}_{t}^{\mu}$ in $L^{2}(\Phi(\mu))$. By the same argument to derive (2.6) from (3.5), we obtain
(3.21) $\|(\tilde{P}_{t_{1}}^{\mu})^{\ast}\|_{L^{\frac{4}{3}}(\Phi(\mu))\rightarrow L^{2}(\Phi(\mu))}=1,\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
So, by Lemma 3.1, we can find constants $c,\lambda>0$ (independent from $K_{I}$), such that, for $K_{I}\in[0,\eta_{0})$,
(3.22) $\mathrm{Ent}((\tilde{P}_{t}^{\mu})^{\ast}\mu_{0}|\Phi(\mu))\leq c\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\Phi(\mu)),\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}).$
It now only remains to repeat the proof of [28, Theorem 3.9]. To be self-contained, we list the main procedure here.
(ii) Reasonning as when we derived (2.8) from (2.7), we deduce from (3.22) that
(3.23) $\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\mu_{0},\Phi(\mu))\leq\tilde{c}\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\mu_{0},\Phi(\mu)),\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d}),$
with $\tilde{c}$ independent from $K_{I}\in[0,\eta_{0})$. By (2.9), it is not difficult to see from Gronwall’s inequality that
(3.24) $\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\Phi(\nu),(\tilde{P}_{t}^{\nu})^{\ast}\Phi(\nu))^{2}\leq\text{\rm{e}}^{(2+2K_{b})t}\int_{0}^{t}K_{I}^{2}\mathbb{W}_{2}(\mu,\nu)^{2}\text{\rm{d}}s,\ \ \mu,\nu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Combining this with (3.23) for $\mu_{0}=\Phi(\nu)$, we have
$\mathbb{W}_{2}(\Phi(\mu),\Phi(\nu))\leq\tilde{c}\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\Phi(\mu),\Phi(\nu))+\text{\rm{e}}^{(1+K_{b})t}\sqrt{t}K_{I}\mathbb{W}_{2}(\mu,\nu),\ \ \mu,\nu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
Let
$\eta_{1}=\left(\inf_{t>\frac{\log{\tilde{c}}}{\lambda}}\frac{\text{\rm{e}}^{(1+K_{b})t}\sqrt{t}}{1-\tilde{c}\text{\rm{e}}^{-\lambda t}}\right)^{-1}.$
When $K_{I}<\min(\eta_{0},\eta_{1})$, the Banach fixed theorem implies that (2.14) has a unique invariant probability measure $\bar{\mu}\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{2d})$.
(Step 2: Exponential ergodicity in $L^{2}$-Wasserstein distance and relative entropy)
Again by (2.9) and Gronwall’s inequality, we get
$\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},(\tilde{P}_{t}^{\bar{\mu}})^{\ast}\mu_{0})^{2}\leq\text{\rm{e}}^{(2+2K_{b})t}\int_{0}^{t}K_{I}^{2}\mathbb{W}_{2}(P_{s}^{\ast}\mu_{0},\bar{\mu})^{2}\text{\rm{d}}s,\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
This together with (3.23) for $\mu=\bar{\mu}$, the triangle inequality and Gronwall’s inequality yields
(3.25) $\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\bar{\mu})^{2}\leq\left(K_{I}^{2}\text{\rm{e}}^{C_{0}t}+2\tilde{c}^{2}\text{\rm{e}}^{-2\lambda t}\right)\mathbb{W}_{2}(\mu_{0},\bar{\mu})^{2},\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$
for some constant $C_{0}>0$ independent from $K_{I}$. This means that we can find $\hat{t}>0$ and $\alpha\in(0,1)$ (independent from $K_{I}$) such that, when $K_{I}$ is small enough,
$\mathbb{W}_{2}(P_{\hat{t}}^{\ast}\mu_{0},\bar{\mu})\leq\alpha\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
By the semigroup property $P_{t}^{\ast}P_{s}^{\ast}=P_{s+t}^{\ast}$ and (3.25), we get (2.15). Next, by (3.4) for $\mu=\bar{\mu}$, the Talagrand inequality holds, i.e.
(3.26) $\mathbb{W}_{2}(\mu_{0},\bar{\mu})^{2}\leq\hat{c}\mathrm{Ent}(\mu_{0}|\bar{\mu}),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),$
for some $\hat{c}>0$ independent from $K_{I}$. Combining the log-Harnack inequality in [29] with (2.15) and (3.26), one may deduce, for some constants $c_{3},c_{4},c_{5}>0$ independent from $K_{I}$,
$\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\bar{\mu})=\mathrm{Ent}(P_{1}^{\ast}P_{t-1}^{\ast}\mu_{0}|P_{1}^{\ast}\bar{\mu})$ $\leq c_{3}\mathbb{W}_{2}(P_{t-1}^{\ast}\mu_{0},\bar{\mu})^{2}$
$\leq c_{4}c\text{\rm{e}}^{-2\lambda(t-1)}\mathbb{W}_{2}(\mu_{0},\bar{\mu})^{2}$
$\leq c_{5}\text{\rm{e}}^{-2\lambda(t-1)}\mathrm{Ent}(\mu_{0}|\bar{\mu}),\ \ t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Hence, the proof is completed. ∎
3.3 Proof of Theorem 2.7
In this section we prove exponential ergodicity for mean field interacting particle systems. First, we reformulate (3.29). For this define
$b^{N}(\mathbf{z})=\left(b\left(z^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{z^{i}}\right)\right)_{1\leq i\leq N},\ \ \mathbf{z}=(z^{1},z^{2},\cdots,z^{N})\in(\mathbb{R}^{2d})^{N}.$
Let $\sigma^{N}=\mathrm{diag}(\sigma_{i})_{1\leq i\leq N}$ with $\sigma_{i}=\sigma$ and $\mathbf{W}_{t}=(W_{t}^{i})_{1\leq i\leq N}$. Then (2.19) can be reformulated as
(3.29) $\left\{\begin{array}[]{ll}\text{\rm{d}}\mathbf{X}_{t}=\mathbf{Y}_{t}\text{\rm{d}}t,\\ \text{\rm{d}}\mathbf{Y}_{t}=b^{N}\left(\mathbf{X}_{t},\mathbf{Y}_{t}\right)\text{\rm{d}}t+\sigma^{N}\text{\rm{d}}\mathbf{W}_{t}.\end{array}\right.$
It is easy to see from (2.9) that
$|b^{N}(\mathbf{z})-b^{N}(\bar{\mathbf{z}})|^{2}$
$=\sum_{i=1}^{N}\left|b\left(z^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{z^{i}}\right)-b\left(\bar{z}^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{\bar{z}^{i}}\right)\right|^{2}$
(3.30) $\leq\sum_{i=1}^{N}\left(2K_{b}^{2}|z_{i}-\bar{z}_{i}|^{2}+2K_{I}^{2}\mathbb{W}_{2}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{z^{i}},\frac{1}{N}\sum_{i=1}^{N}\delta_{\bar{z}^{i}}\right)^{2}\right)$
$\leq\sum_{i=1}^{N}\left(2K_{b}^{2}|z_{i}-\bar{z}_{i}|^{2}+2K_{I}^{2}\frac{1}{N}\sum_{i=1}^{N}|z^{i}-\bar{z}^{i}|^{2}\right)$
$=(2K_{b}^{2}+2K_{I}^{2})\sum_{i=1}^{N}|z^{i}-\bar{z}^{i}|^{2},\ \ \mathbf{z}=(z^{1},z^{2},\cdots,z^{N}),\bar{\mathbf{z}}=(\bar{z}^{1},\bar{z}^{2},\cdots,\bar{z}^{N})\in(\mathbb{R}^{2d})^{N}.$
This implies that under (2.9), (2.19) is well-posed and we denote by $P^{N}_{t}$ the associated semigroup. Let $(P_{t}^{N})^{\ast}\nu^{N}$ be the distribution of the solution to (2.19) with initial distribution $\nu^{N}\in\mathscr{P}((\mathbb{R}^{2d})^{N})$.
Proof of Theorem 2.7.
Reasoning as in the proof of Theorem 2.6, unless explicitly mentioned otherwise, constants appearing in the proofs are all by default independent from $K_{I}$ small enough (and, of course, from $N$).
(Step 1: $N$-particle partial dissipativity) Let $\mathbf{x}=(x^{i})_{1\leq i\leq N}$, $\mathbf{y}=(y^{i})_{1\leq i\leq N}$, $\bar{\mathbf{x}}=(\bar{x}^{i})_{1\leq i\leq N}$, $\bar{\mathbf{y}}=(\bar{y}^{i})_{1\leq i\leq N}\in(\mathbb{R}^{d})^{N}$. $\mathbf{z}=(\mathbf{x},\mathbf{y})$, $\bar{\mathbf{z}}=(\bar{\mathbf{x}},\bar{\mathbf{y}})$. Firstly, it follows from ((C)) and (2.9) that
$\langle r^{2}(\mathbf{x}-\bar{\mathbf{x}})+rr_{0}(\mathbf{y}-\bar{\mathbf{y}}),\mathbf{y}-\bar{\mathbf{y}}\rangle$
$+\langle(\mathbf{y}-\bar{\mathbf{y}})+rr_{0}(\mathbf{x}-\bar{\mathbf{x}}),b^{N}(\mathbf{x},\mathbf{y})-b^{N}(\bar{\mathbf{x}},\bar{\mathbf{y}})\rangle$
$=\sum_{i=1}^{N}\left\langle r^{2}(x^{i}-\bar{x}^{i})+rr_{0}(y^{i}-\bar{y}^{i}),y^{i}-\bar{y}^{i}\right\rangle$
(3.31) $+\sum_{i=1}^{N}\langle(y^{i}-\bar{y}^{i})+rr_{0}(x^{i}-\bar{x}^{i}),b(x^{i},y^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{(x^{i},y^{i})})-b(\bar{x}^{i},\bar{y}^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{(x^{i},y^{i})})\rangle$
$+\sum_{i=1}^{N}\langle(y^{i}-\bar{y}^{i})+rr_{0}(x^{i}-\bar{x}^{i}),b(\bar{x}^{i},\bar{y}^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{(x^{i},y^{i})})-b(\bar{x}^{i},\bar{y}^{i},\frac{1}{N}\sum_{i=1}^{N}\delta_{(\bar{x}^{i},\bar{y}^{i})})\rangle$
$\leq\sum_{i=1}^{N}(-\theta+K_{I}(1+|rr_{0}|))(|x^{i}-\bar{x}^{i}|^{2}+|y^{i}-\bar{y}^{i}|^{2})$
$+\sum_{i=1}^{N}(C_{0}+\theta)(|x^{i}-\bar{x}^{i}|^{2}+|y^{i}-\bar{y}^{i}|^{2})1_{\{|x^{i}-\bar{x}^{i}|^{2}+|y^{i}-\bar{y}^{i}|^{2}<R^{2}\}}$
$\leq(-\theta+K_{I}(1+|rr_{0}|))|\mathbf{z}-\bar{\mathbf{z}}|^{2}+(C_{0}+\theta)|\mathbf{z}-\bar{\mathbf{z}}|^{2}1_{\{|\mathbf{z}-\bar{\mathbf{z}}|^{2}<NR^{2}\}}.$
So, when $K_{I}$ is small enough (independently from $N$), it follows from [28, Proof of Theorem 3.3] that (2.19) has a unique invariant probability measure $\bar{\mu}^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N})$.
(Step 2: introducing the non-linear equilibrium) Assuming that $K_{I}$ is small enough so that the conclusion of Theorem 2.6 holds, let $\bar{\mu}$ be the unique invariant probability measure of the McKean-Vlasov SDE (2.22) and $\tilde{P}_{t}^{\bar{\mu}}$ be the semigroup associated to the decoupled SDE
(3.34) $\left\{\begin{array}[]{ll}\text{\rm{d}}\tilde{X}_{t}^{\bar{\mu}}=\tilde{Y}_{t}^{\bar{\mu}}\text{\rm{d}}t,\\ \text{\rm{d}}\tilde{Y}_{t}^{\bar{\mu}}=b(\tilde{X}_{t}^{\bar{\mu}},\tilde{Y}_{t}^{\bar{\mu}},\bar{\mu})\text{\rm{d}}t+\sigma\text{\rm{d}}W_{t}.\end{array}\right.$
Denote $\tilde{X}_{t}^{i,z,\bar{\mu}}$ the solution to (3.34) with $W_{t}^{i}$ replacing $W_{t}$ and initial value $z\in\mathbb{R}^{2d}$. For any $\mathbf{z}=(z^{1},z^{2},\cdots,z^{N}),F\in\mathscr{B}_{b}((\mathbb{R}^{2d})^{N})$, define
$(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N}(F)(\mathbf{z})=\mathbb{E}F(\tilde{X}_{t}^{1,z^{1},\bar{\mu}},\tilde{X}_{t}^{2,z^{2},\bar{\mu}},\cdots,\tilde{X}_{t}^{N,z^{N},\bar{\mu}}),$
and for any $\nu^{N}\in\mathscr{P}((\mathbb{R}^{2d})^{N})$, define
$(\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N})(F)=\int_{(\mathbb{R}^{2d})^{N}}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N}(F)\text{\rm{d}}\nu^{N},\ \ F\in\mathscr{B}_{b}((\mathbb{R}^{2d})^{N}).$
By the triangle inequality, we have
(3.35) $\mathbb{W}_{2}((P_{t}^{N})^{\ast}\nu^{N},\bar{\mu}^{\otimes N})$ $\leq\mathbb{W}_{2}((P_{t}^{N})^{\ast}\nu^{N},\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N})+\mathbb{W}_{2}(\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N},\bar{\mu}^{\otimes N})$
$=:I_{1}+I_{2}.$
Next, we estimate $I_{1},I_{2}$ one after the other in the two next steps.
(Step 3: contraction for $(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N}$) We start with $I_{2}$. Firstly, it follows from (3.23) for $\mu=\bar{\mu}$ that
$\mathbb{W}_{2}((\tilde{P}_{t}^{\bar{\mu}})^{\ast}\nu,\bar{\mu})\leq c_{0}\text{\rm{e}}^{-\lambda_{0}t}\mathbb{W}_{2}(\nu,\bar{\mu}),\ \ \nu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
By (3.21) for $\mu=\bar{\mu}$, we have
$\|(\tilde{P}_{t_{1}}^{\bar{\mu}})^{\ast}\|_{L^{\frac{4}{3}}(\bar{\mu})\rightarrow L^{2}(\bar{\mu})}=1.$
We first claim the tensor property of hypercontractivity, i.e.
(3.36) $\|[(\tilde{P}_{t_{1}}^{\bar{\mu}})^{\ast}]^{\otimes N}\|_{L^{\frac{4}{3}}(\bar{\mu}^{\otimes N})\rightarrow L^{2}(\bar{\mu}^{\otimes N})}=1.$
By the method of induction, it is sufficient to prove the case $N=2$. Indeed, for any $g\in L^{\frac{4}{3}}(\bar{\mu}^{\otimes 2})$, let $g_{y}(z)=(\tilde{P}_{t}^{\bar{\mu}})^{\ast}g(z,\cdot)(y)$. Then it holds
$|[(\tilde{P}_{t}^{\bar{\mu}})^{\ast}]^{\otimes 2}\|_{L^{\frac{4}{3}}(\bar{\mu}^{\otimes 2})\rightarrow L^{2}(\bar{\mu}^{\otimes 2})}^{2}$
$=\sup_{\|g\|_{L^{\frac{4}{3}}(\bar{\mu}^{\otimes 2})}\leq 1}\int_{\mathbb{R}^{2d}}\int_{\mathbb{R}^{2d}}|[(\tilde{P}_{t}^{\bar{\mu}})^{\ast}g_{y}](x)|^{2}\mu(\text{\rm{d}}x)\mu(\text{\rm{d}}y)$
$\leq\sup_{\|g\|_{L^{\frac{4}{3}}(\bar{\mu}^{\otimes 2})}\leq 1}\int_{\mathbb{R}^{2d}}\left(\int_{\mathbb{R}^{2d}}|(\tilde{P}_{t}^{\bar{\mu}})^{\ast}g(x,\cdot)(y)|^{\frac{4}{3}}\mu(\text{\rm{d}}x)\right)^{\frac{3}{2}}\mu(\text{\rm{d}}y)$
$\leq\sup_{\|g\|_{L^{\frac{4}{3}}(\bar{\mu}^{\otimes 2})}\leq 1}\left(\int_{\mathbb{R}^{2d}}\left(\int_{\mathbb{R}^{2d}}|(\tilde{P}_{t}^{\bar{\mu}})^{\ast}g(x,\cdot)(y)|^{2}\mu(\text{\rm{d}}y)\right)^{\frac{2}{3}}\mu(\text{\rm{d}}x)\right)^{\frac{3}{2}}$
$\leq\sup_{\|g\|_{L^{\frac{4}{3}}(\bar{\mu}^{\otimes 2})}\leq 1}\left(\int_{\mathbb{R}^{2d}}\int_{\mathbb{R}^{2d}}g(x,y)|^{\frac{4}{3}}\mu(\text{\rm{d}}y)\mu(\text{\rm{d}}x)\right)^{\frac{3}{2}}$
$\leq 1,$
where in the second inequality, we use Minkowski’s inequality. Then by Lemma 3.1 and (3.36), we can find some constants $c_{0},\lambda_{0}>0$ independent of $N$ such that
(3.37) $\mathrm{Ent}(\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N}|\bar{\mu}^{\otimes N})\leq c_{0}\text{\rm{e}}^{-\lambda_{0}t}\mathrm{Ent}(\nu^{N}|\bar{\mu}^{\otimes N}),\ \ \nu^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N}).$
Moreover, by (3.4) for $\mu=\bar{\mu}$, we have
(3.38) $\bar{\mu}(f\log f)\leq C_{\mathrm{LS}}\bar{\mu}(|\nabla\sqrt{f}|^{2}),\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),f>0,\bar{\mu}(f)=1.$
This together with the tensor property of log-Sobolev inequality implies
(3.39) $\bar{\mu}^{\otimes N}(f\log f)\leq C_{\mathrm{LS}}\bar{\mu}^{\otimes N}(|\nabla\sqrt{f}|^{2}),\ \ f\in C_{0}^{\infty}((\mathbb{R}^{2d})^{N}),f>0,\bar{\mu}^{\otimes N}(f)=1,$
which itself implies the Talangrand inequality
(3.40) $\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N})^{2}\leq 4C_{\mathrm{LS}}\mathrm{Ent}(\nu^{N}|\bar{\mu}^{\otimes N}),\ \ \nu^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N}).$
By [53], the log-Harnack inequality holds:
(3.41) $\tilde{P}_{1}^{\bar{\mu}}\log f(x)\leq\log\tilde{P}_{1}^{\bar{\mu}}f(\bar{x})+c_{1}|x-\bar{x}|^{2}.$
It is easy to see from this that
(3.42) $(\tilde{P}_{1}^{\bar{\mu}})^{\otimes N}\log f(\mathbf{x})\leq\log(\tilde{P}_{1}^{\bar{\mu}})^{\otimes N}f(\bar{\mathbf{x}})+c_{1}|\mathbf{x}-\bar{\mathbf{x}}|^{2}.$
From this, [61, Proposition 1.4.4(3)] gives
(3.43) $\mathrm{Ent}(\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N}|\bar{\nu}^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N})\leq c_{1}\mathbb{W}_{2}(\nu^{N},\bar{\nu}^{N})^{2}.$
Then by the same argument to derive (2.8) from (2.7), we gain from (3.37) that
(3.44) $I_{2}=\mathbb{W}_{2}(\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N},\bar{\mu}^{\otimes N})\leq c_{0}\text{\rm{e}}^{-\lambda_{0}t}\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N}),\ \ \nu^{N}\in\mathscr{P}_{2}((\mathbb{R}^{2d})^{N}).$
(Step 4: synchronous coupling) In order to estimate $I_{1}$, we will use a synchronous coupling. Consider
$\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}^{i,N}=Y_{t}^{i,N}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}^{i,N}=b(X_{t}^{i,N},Y_{t}^{i,N},\frac{1}{N}\sum_{i=1}^{N}\delta_{(X_{t}^{i,N},Y_{t}^{i,N})})\text{\rm{d}}t+\sigma\text{\rm{d}}W_{t}^{i},\ \ \mathscr{L}_{(X_{0}^{i,N},Y_{0}^{i,N})_{1\leq i\leq N}}=\nu^{N},\end{array}\right.$
and
$\left\{\begin{array}[]{ll}\text{\rm{d}}\bar{X}_{t}^{i,N}=\bar{Y}_{t}^{i,N}\text{\rm{d}}t,\\ \text{\rm{d}}\bar{Y}_{t}^{i,N}=b(\bar{X}_{t}^{i,N},\bar{Y}_{t}^{i,N},\bar{\mu})\text{\rm{d}}t+\sigma\text{\rm{d}}W_{t}^{i},\ \ (\bar{X}_{0}^{i,N},\bar{Y}_{0}^{i,N})_{1\leq i\leq N}=(X_{0}^{i,N},Y_{0}^{i,N})_{1\leq i\leq N},\end{array}\right.$
Then it holds $I_{1}^{2}\leq\mathbb{E}\sum_{i=1}^{N}(|X_{t}^{i,N}-\bar{X}_{t}^{i,N}|^{2}+|Y_{t}^{i,N}-\bar{Y}_{t}^{i,N}|^{2})$. By (2.9), we have
$\sum_{i=1}^{N}(|X_{t}^{i,N}-\bar{X}_{t}^{i,N}|^{2}+|Y_{t}^{i,N}-\bar{Y}_{t}^{i,N}|^{2})$
$\leq C_{0}\int_{0}^{t}\sum_{i=1}^{N}(|X_{s}^{i,N}-\bar{X}_{s}^{i,N}|^{2}+|Y_{s}^{i,N}-\bar{Y}_{s}^{i,N}|^{2})\text{\rm{d}}s$
$+K_{I}^{2}N\int_{0}^{t}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(X_{s}^{i,N},Y_{s}^{i,N})},\bar{\mu})^{2}\text{\rm{d}}s$
$\leq C_{0}\int_{0}^{t}\sum_{i=1}^{N}(|X_{s}^{i,N}-\bar{X}_{s}^{i,N}|^{2}+|Y_{s}^{i,N}-\bar{Y}_{s}^{i,N}|^{2})\text{\rm{d}}s$
$+2K_{I}^{2}N\int_{0}^{t}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(X_{s}^{i,N},Y_{s}^{i,N})},\frac{1}{N}\sum_{i=1}^{N}\delta_{(\bar{X}_{s}^{i,N},\bar{Y}_{s}^{i,N})})^{2}\text{\rm{d}}s$
$+2K_{I}^{2}N\int_{0}^{t}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\bar{X}_{s}^{i,N},\bar{Y}_{s}^{i,N})},\bar{\mu})^{2}\text{\rm{d}}s$
$\leq(C_{0}+2K_{I}^{2})\int_{0}^{t}\sum_{i=1}^{N}(|X_{s}^{i,N}-\bar{X}_{s}^{i,N}|^{2}+|Y_{s}^{i,N}-\bar{Y}_{s}^{i,N}|^{2})\text{\rm{d}}s$
$+2K_{I}^{2}N\int_{0}^{t}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\bar{X}_{s}^{i,N},\bar{Y}_{s}^{i,N})},\bar{\mu})^{2}\text{\rm{d}}s.$
For any $\pi^{N}\in\mathscr{C}(\nu^{N}(\tilde{P}_{s}^{\bar{\mu}})^{\otimes N},\bar{\mu}^{\otimes N})$, it holds
$\mathbb{E}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\bar{X}_{s}^{i,N},\bar{Y}_{s}^{i,N})},\bar{\mu})^{2}$
$=\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(x^{i},y^{i})},\bar{\mu})^{2}\text{\rm{d}}(\nu^{N}(P_{s}^{\bar{\mu}})^{\otimes N})$
$=\int_{(\mathbb{R}^{2d})^{N}\times(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(x^{i},y^{i})},\bar{\mu})^{2}\text{\rm{d}}\pi^{N}$
$\leq 2\int_{(\mathbb{R}^{2d})^{N}\times(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(x^{i},y^{i})},\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})})^{2}\text{\rm{d}}\pi^{N}$
$+2\int_{(\mathbb{R}^{2d})^{N}\times(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\pi^{N}$
$\leq 2\frac{1}{N}\int_{(\mathbb{R}^{2d})^{N}\times(\mathbb{R}^{2d})^{N}}\sum_{i=1}^{N}(|x^{i}-\tilde{x}^{i}|^{2}+|y^{i}-\tilde{y}^{i}|^{2})\text{\rm{d}}\pi^{N}$
$+2\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\bar{\mu}^{\otimes N}.$
Taking infimum in $\pi^{N}\in\mathscr{C}(\nu^{N}(\tilde{P}_{s}^{\bar{\mu}})^{\otimes N},\bar{\mu}^{\otimes N})$, we conclude from (3.44) that
$\mathbb{E}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\bar{X}_{s}^{i,N},\bar{Y}_{s}^{i,N})},\bar{\mu})^{2}$ $\leq 2\frac{1}{N}\mathbb{W}_{2}(\nu^{N}(\tilde{P}_{s}^{\bar{\mu}})^{\otimes N},\bar{\mu}^{\otimes N})^{2}$
$+2\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\bar{\mu}^{\otimes N}$
$\leq 2c_{0}^{2}\text{\rm{e}}^{-2\lambda_{0}s}\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N})^{2}\frac{1}{N}$
$+2\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\bar{\mu}^{\otimes N}.$
Taking expectation and using Gronwall’ s inequality, we deduce
$I_{1}^{2}$ $=\mathbb{W}_{2}((P_{t}^{N})^{\ast}\nu^{N},\nu^{N}(\tilde{P}_{t}^{\bar{\mu}})^{\otimes N})^{2}\leq\mathbb{E}\sum_{i=1}^{N}(|X_{t}^{i,N}-\bar{X}_{t}^{i,N}|^{2}+|Y_{t}^{i,N}-\bar{Y}_{t}^{i,N}|^{2})$
(3.45) $\leq 4\text{\rm{e}}^{(C_{0}+2K_{I}^{2})t}\int_{0}^{t}K_{I}^{2}c_{0}^{2}\text{\rm{e}}^{-2\lambda_{0}s}\text{\rm{d}}s\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N})^{2}$
$+\text{\rm{e}}^{(C_{0}+2K_{I}^{2})t}4K_{I}^{2}N\int_{0}^{t}\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\bar{\mu}^{\otimes N}\text{\rm{d}}s.$
(Step 5: combining $I_{1}$ and $I_{2}$) Plugging (3.44) and (3.3) in (3.35) gives
$\mathbb{W}_{2}((P_{t}^{N})^{\ast}\nu^{N},\bar{\mu}^{\otimes N})^{2}$ $\leq 2c_{0}^{2}\text{\rm{e}}^{-2\lambda_{0}t}\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N})^{2}$
(3.46) $+8\text{\rm{e}}^{(C_{0}+2K_{I}^{2})t}\int_{0}^{t}K_{I}^{2}c_{0}^{2}\text{\rm{e}}^{-2\lambda_{0}s}\text{\rm{d}}s\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N})^{2}$
$+2\text{\rm{e}}^{(C_{0}+2K_{I}^{2})t}4K_{I}^{2}N\int_{0}^{t}\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\bar{\mu}^{\otimes N}\text{\rm{d}}s.$
In view of [28, (3.12)], there exists a constant $\varepsilon>0$ such that
$\bar{\mu}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})<\infty.$
This together with [20, Theorem 1] implies
$\int_{(\mathbb{R}^{2d})^{N}}\mathbb{W}_{2}(\frac{1}{N}\sum_{i=1}^{N}\delta_{(\tilde{x}^{i},\tilde{y}^{i})},\bar{\mu})^{2}\text{\rm{d}}\bar{\mu}^{\otimes N}\leq C_{2d}R_{d}(N).$
Using this in (3.3), we see that we can find constants $\hat{t},\bar{c},K_{*}>0$ and $\alpha\in(0,\frac{1}{9})$ such that, when $K_{I}<K_{*}$,
(3.47) $\mathbb{W}_{2}((P_{\hat{t}}^{N})^{\ast}\nu^{N},\bar{\mu}^{\otimes N})^{2}\leq\alpha\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{\otimes N})^{2}+\bar{c}NR_{d}(N).$
(Step 6: conclusion) Taking $\nu^{N}=\bar{\mu}^{N}$ in (3.47) gives
$\mathbb{W}_{2}(\bar{\mu}^{N},\bar{\mu}^{\otimes N})^{2}\leq\frac{\bar{c}NR_{d}(N)}{1-\alpha}.$
So, we get
$\mathbb{W}_{2}((P_{\hat{t}}^{N})^{\ast}\nu^{N},\bar{\mu}^{\otimes N})^{2}\leq 2\alpha\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{N})^{2}+\frac{2\alpha\bar{c}NR_{d}(N)}{1-\alpha}.$
Note that
$\mathbb{W}_{2}((P_{\hat{t}}^{N})^{\ast}\nu^{N},\bar{\mu}^{N})^{2}\leq 2\mathbb{W}_{2}((P_{\hat{t}}^{N})^{\ast}\nu^{N},\bar{\mu}^{\otimes N})^{2}+2\mathbb{W}_{2}(\bar{\mu}^{N},\bar{\mu}^{\otimes N})^{2}.$
We conclude that
(3.48) $\mathbb{W}_{2}((P_{\hat{t}}^{N})^{\ast}\nu^{N},\bar{\mu}^{N})^{2}\leq 4\alpha\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{N})^{2}+\frac{(4\alpha\bar{c}+2\bar{c})NR_{d}(N)}{1-\alpha}.$
Moreover, it follows from (3.3) that
$\sup_{t\in[0,\hat{t}]}\mathbb{W}_{2}((P_{t}^{N})^{\ast}\nu^{N},\bar{\mu}^{N})^{2}\leq\tilde{c}\mathbb{W}_{2}(\nu^{N},\bar{\mu}^{N})^{2}+\tilde{c}NR_{d}(N)$
This combined with (3.48) as well as the semigroup property implies (2.7). Finally, by (3.3), the log-Harnack inequality
$\mathrm{Ent}((P_{t}^{N})^{\ast}\nu^{N}|(P_{t}^{N})^{\ast}\bar{\nu}^{N})\leq ct^{-3}\mathbb{W}_{2}(\nu^{N},\bar{\nu}^{N})^{2},\ \ t\in(0,1]$
holds for some constant $c$ independent of $N$, see for instance [53]. Applying this inequality for $\bar{\nu}^{N}=\bar{\mu}^{N}$ and combining with (2.7) and the semigroup property, we obtain (2.23) and the proof is completed. ∎